# Experimental single-copy distillation of quantumness from higher-dimensional entanglement

### v1

@@ Abstract
Introduction.–Ever since its discovery, quantum entanglement Horodecki et al. (2009) has played a fascinating role in shaping Bell (1964) our understanding of the physical world. In the modern era, entanglement takes the role of a powerful resource, enabling quantum computation Jozsa and Linden (2003); Vidal (2003), quantum-enhanced communications Ekert (1991); Bennett and Wiesner (1992), and quantum metrology Tóth and Apellaniz (2014), etc. To take advantage of these possibilities at a large scale, a quantum internet Wehner et al. (2018) where we can readily perform quantum communications between any two points is clearly desirable. Accordingly, the possibility of distributing entanglement in one way or another (see, e.g., Żukowski et al. (1993); Yin et al. (2017)) will be more than welcome.
@@
Even with ideal preparation, the quality of entanglement can easily degrade over time, either during the storage or distribution stage. Consequently, much effort has been devoted to understanding how one can recover maximally entangled Bell pairs via entanglement distillation or purification protocols Bennett et al. (1996); Deutsch et al. (1996). Indeed, for quantum states violating the reduction criterion Horodecki and Horodecki (1999), iterative single-copy local filtering operations (ScLFs) Gisin (1996) followed by collective local operations Bennett et al. (1996); Horodecki and Horodecki (1999) distill, with a nonzero probability, quantum states having an arbitrarily high Bell-state fidelity.
@@
Nonetheless, some bipartite entangled states, e.g., those having positive-partial-transposition Peres (1996) (PPT), are undistillable Horodecki et al. (1998). There is even evidence DiVincenzo et al. (2000); Dür et al. (2000) suggesting that some non-PPT (NPPT) quantum states, such as the Werner states Werner (1989), are undistillable. Since all two-qubit entangled states are distillable Horodecki et al. (1997), this intriguing phenomenon of bound entanglement Horodecki et al. (1998) only occurs among higher-dimensional (HD) mixed entangled states. In contrast, pure HD entanglement may offer an advantage over its qubit counterparts in entanglement distribution Steinlechner et al. (2017); Ecker et al. (2019); Hu et al. (2020a) and various information-processing tasks (see, e.g., Liu et al. (2002); Liang et al. (2003); Vértesi et al. (2010) for some theoretical proposals).
@@
To date, it remains unknown Horodecki et al. (2022) whether all NPPT entangled states are distillable. Even among those that are distillable, the number of copies required may be arbitrarily large Watrous (2004); Fang and Liu (2020, 2022); Regula (2022). Moreover, to distill via the (generalized) recurrence protocol Bennett et al. (1996); Horodecki and Horodecki (1999), the input state to the protocol must already have a large enough fully-entangled fraction Horodecki et al. (1999) (to the point that it even ensures its usefulness Popescu (1994) for teleportation). To meet this requirement, the initial noisy entangled state may thus have to go through additional ScLFs before being subjected to the collective distillation operations. Even though latter, more technically challenging, operations have been achieved for two copies in various experiments Pan et al. (2001, 2003); Zhao et al. (2003); Yamamoto et al. (2003); Walther et al. (2005); Reichle et al. (2006); Kalb et al. (2017) (see also Hu et al. (2021a); Ecker et al. (2021)), we are still far from recovering a near-to-perfect Bell pair via conventional distillation schemes.
@@
In contrast, for the sake of recovering quantumness or quantum advantage over classical resources in information processing tasks, ScLFs often suffice. For instance, they can help Zhao et al. (2017) distinguish different types of entanglement, recover usefulness for secure communication Singh et al. (2021), unveil hidden Pramanik et al. (2019); Nery et al. (2020); Hao et al. (2024); Zhang et al. (2024) quantum steerability Wiseman et al. (2007); Uola et al. (2020), hidden Li et al. (2021) teleportation power Popescu (1994) (see also Masanes (2006)), as well as the (seemingly) hidden Bell-nonlocality of certain two-qubit states Gisin (1996); Kwiat et al. (2001); Wang et al. (2006, 2020) and HD states Popescu (1995); Kumari (2024) (see also Masanes et al. (2008); Liang et al. (2012)). Here, building on the results from Popescu (1995); Liang et al. (2012); Li et al. (2021), we experimentally demonstrate the activation of various desired properties via the same ScLF for one classic family of HD quantum states—the two-qutrit Werner states. Although tremendous efforts have been devoted to the generation and detection of HD entanglement and their nonlocal properties (see, e.g., Thew et al. (2004); Huber et al. (2010); Schwarz et al. (2016); Martin et al. (2017); Islam et al. (2017); Bavaresco et al. (2018); Zeng et al. (2018); Wang et al. (2018); Hu et al. (2018); Lu et al. (2020); Dada et al. (2011); Hu et al. (2020b, 2021b); Designolle et al. (2021); Qu et al. (2022a, b)), our work gives the first experimental demonstration of the single-copy distillation of quantumness from HD entanglement.
@@
where $\mathbb{I}_{d^{2}}$ is the identity operator acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$, $n_{\pm}^{d}:={\rm tr}\Pi_{\pm}^{d}=\frac{d(d\pm 1)}{2}$ and $V$ is the swap operator, i.e., $V|\psi\rangle_{A}|\phi\rangle_{B}=|\phi\rangle_{A}|\psi\rangle_{B}$ for all $|\psi\rangle,|\phi\rangle\in\mathbb{C}^{d}$. $W^{(d)}(v)$ are PPT Dür et al. (2000) and separable Werner (1989) if and only if (iff ) $v\geq\frac{1}{2}$. However, entangled Werner states are only known to be distillable when Dür et al. (2000); DiVincenzo et al. (2000) $v<v_{\text{distill}}\equiv\frac{d+1}{4d-2}$. In this regard, a state $\rho$ is said to be distillable Horodecki and Horodecki (2001) if it is $k$-distillable for some natural number $k$, i.e., if there exists qubit projectors $\Pi_{A},\Pi_{B}$ such that $\frac{\Pi_{A}\otimes\Pi_{B}\rho^{\otimes k}\Pi_{A}\otimes\Pi_{B}}{{\rm tr}(\Pi_{A}\otimes\Pi_{B}\rho^{\otimes k}\Pi_{A}\otimes\Pi_{B})}$ is entangled. To determine the existence of NPPT bound entangled states Horodecki et al. (2022), it suffices Dür et al. (2000) to consider $W^{(d)}(v)$ for $d>2$.
@@
HD Werner states are peculiar, however, not only because of their distillability properties but also because they are useless in various tasks. Indeed, all HD $W^{(d)}(v)$’s satisfy the reduction criterion (RC) of separability Horodecki and Horodecki (1999). Consequently, even those that are distillable cannot be fed directly to the generalized recurrence protocol Horodecki and Horodecki (1999) for entanglement distillation. Their compliance with RC also renders them useless for quantum teleportation Horodecki and Horodecki (1999) and superdense coding Bruß et al. (2004). However, as shown in Li et al. (2021), an appropriate ScLF (in the form of qubit projection) can activate their usefulness for teleportation whenever $v<v_{\text{distill}}$.
@@
Furthermore, Werner states with $v\geq v_{\text{steer}}\equiv\frac{d+1}{2d^{2}}$ are the first known examples of entangled quantum states not violating any Bell Werner (1989) or steering Wiseman et al. (2007) inequality with projective measurements (see Fig. 1). However, Popescu Popescu (1995) showed that after a successful qubit ScLF, it is possible for $W^{(d\geq 5)}(v_{\text{steer}})$ to violate the Clauser-Horne-Shimony-Holt (CHSH) Bell inequality Clauser et al. (1969). More generally, $W^{(d)}(v)$’s are known (Johnson and Viola, 2013, Theorem 6) to admit a symmetric $(1,k)$ or $(k,1)$-extension Terhal et al. (2003); Doherty et al. (2002) iff $v\geq v_{\text{Sym}}\equiv\frac{1}{2}\left(1-\frac{d-1}{k}\right)$. In other words, for these states, there exists a $(1+k)$-partite qudit state $\rho$ that recovers $W^{(d)}(v)$ after tracing out any $(k-1)$ copies of Alice’s (Bob’s) subsystems. It then follows Terhal et al. (2003) that $W^{(d)}(v)$ for $v\geq v_{\text{Sym}}$ cannot violate any Bell inequality with $\min\{s_{a},s_{b}\}\leq k$ where $s_{a}$ ($s_{b}$) is the number of measurement settings for Alice (Bob). [^1] To construct a local-hidden-variable model (LHVM) for $W^{(d)}(v)$, it suffices to seek for a $(1,k)$-symmetric quasiextension Terhal et al. (2003) where $\rho$ is even an entanglement witness, i.e., not being positive semidefinite. However, our computation results suggest that this relaxation does not lead to the construction of LHVM for a wider range of $W^{(d)}(v)$. On the other hand, our findings also suggest that $W^{(d)}(v)$ admit a $(k,1)$ bosonic symmetric extension Doherty (2014) iff $W^{(d=2)}(v)$ does. See Appendix B.1.1 for details. In particular, since $v_{\text{Sym}}$ vanishes when $k=d-1$, $W^{(d)}(v)$ satisfy Terhal et al. (2003) all Bell inequalities with $d-1$ or fewer measurement bases at one side. For example, all two-qutrit Werner states cannot violate any two-or-fewer-setting Bell inequality on one side. Again, a qubit ScLF reveals Liang et al. (2012) the Bell-nonlocality (possibly) hidden in some of these states, which we demonstrate experimentally in this work.
@@
How does a qubit ScLF enable the distillation of many of these useful properties? We can gain some intuition into this by first noting that $\Pi_{+}^{d}+\Pi_{-}^{d}=\mathbb{I}_{d^{2}}$ and Popescu (1995) $\Pi_{-}^{d}=\sum_{i<j}\varrho_{0}^{i,j}$ where $\varrho_{0}^{i,j}=|\Psi^{-}_{i,j}\rangle\!\langle\Psi^{-}_{i,j}|$ is a singlet projector defined on the two-qubit subspace $\mathcal{H}_{i,j}:=\text{span}\{|k\ell\rangle:=|k\rangle|\ell\rangle\}_{k,\ell=i,j}$. Then, by defining $p:=\frac{1}{d}$, $q:=1-\frac{2d}{d+1}v$, and noting that $n_{-}^{d}$ is also the binomial coefficient of $d$ chooses 2, we can rewrite $W^{(d)}(v)$ as a uniform mixture of two-qubit states $\varrho^{i,j}$:
@@
where $\varrho_{1}^{i,j}=\frac{|ii\rangle\!\langle ii|+|jj\rangle\!\langle jj|}{2}$, $\varrho_{2}^{i,j}=\frac{|ij\rangle\!\langle ij|+|ji\rangle\!\langle ji|}{2}$. In this decomposition, only the two-qubit singlet projectors $\{\varrho_{0}^{i,j}\}_{i,j}$ contribute to the entanglement of $W^{(d)}(v)$. The act of a qubit projection, say, $\Pi_{2}:=|0\rangle\!\langle 0|+|1\rangle\!\langle 1|$, on both sides then keeps the maximally entangled $\varrho_{0}^{0,1}$ and the separable $\frac{\mathbb{I}_{4}}{4}$ with some probability while getting rid of the remaining components in all other two-qubit subspaces, thereby effectively “purifying” the entanglement contained in $W^{(d)}(v)$. From Eq. 2a, we see that upon a successful qubit projection, since $\Pi_{2}\otimes\Pi_{2}W^{(d)}(v)\Pi_{2}\otimes\Pi_{2}\propto W^{(d=2)}(v^{\prime})$, we obtain a two-qubit Werner state with a symmetric weight $v^{\prime}=\frac{3(d-1)v}{(d+1)(1-v)+3(d-1)v}$. Note that $v^{\prime}\geq v$ for all $d>2$.
@@
Experimental Werner states and filtering operation.–The decomposition of Eq. 2 not only sheds light on the relevance of a qubit ScLF but also provides a systematic way for preparing an entangled $W^{(d)}(v)$ for an arbitrary $d$ via generating and mixing the various two-qubit states $\varrho^{i,j}(q)$ residing in $\mathcal{H}_{i,j}$. To this end, we start by preparing a pair of maximally entangled photons in the polarization degree of freedom (DOF) $|\Psi^{+}\rangle_{\text{AB}}=\frac{1}{\sqrt{2}}(|H\rangle|V\rangle+|V\rangle|H\rangle)_{\text{AB}}$, where $H$ and $V$ denote, respectively, horizontal and vertical polarization. As shown in Fig. 2a, the entangled photons are generated via the spontaneous parameter down conversion process (SPDC) from a periodically-poled potassium titanyl phosphate (PPKTP) crystal in a Sagnac interferometer, which is bidirectionally pumped by a laser with central wavelength at 405 nm. By setting a HWP at 0 on photon B, $|\Psi^{+}\rangle_{\text{AB}}$ is transformed into $\varrho_{0}^{H,V}=|\Psi^{-}\rangle_{\text{AB}}\langle\Psi^{-}|$. Then, we use two BSs, two PBSs and two tunable LAAs (with transmittance of $q$ and $1-q$, respectively) to partially decohere $\varrho_{0}^{H,V}$ and prepare the mixed state $\varrho^{H,V}(q)$ in Eq. 2 by randomly rotating a HWP between $45^{\circ}$ and $0^{\circ}$. Finally, we obtain the $d$-dimensional Werner states $W^{(d)}(v)$ from $\varrho^{H,V}(q)$ with a qubit-to-qudit mapping, where the qudit is encoded in the path DOF $\{|l_{k}\rangle\}$ and the mapping is achieved with two optical networks each consisting of $d-1$ beam displacers (BDs) and a series of HWPs on photon A and B. For $d=3$ in this work, we use two BDs and two HWPs at each side (as shown in Fig. 2a) to prepare eleven $W^{(d=3)}(v)$ with $v\in[0,0.5]$ in steps of 0.05. See Appendix A.2 for further details.
@@
ScLF and distillation of quantumness.–With our encoding of the quantum information, a qubit filtering—relevant for distilling the various quantum features alluded to above—is straightforward. As shown in Fig. 2b, to keep only contributions to the two lower optical paths $\{|l_{1}\rangle,|l_{2}\rangle\}$, we block the light propagating through the upper optical path $|l_{0}\rangle$. After ScLF, the two-qubit state $\rho_{\text{expt},f}(v)$ encoded in $\{|l_{1}\rangle,|l_{2}\rangle\}$ is mapped back to the polarization DOF and then reconstructed using a HWP, a QWP and a PBS, and their fidelity ranges roughly from 0.934 to 0.995, cf. Table 4. In the following, we compare the properties of $\rho_{\text{expt}}(v)$ and $\rho_{\text{expt},f}(v)$ to illustrate how ScLF has enabled the recovery of quantum features missing in $\rho_{\text{expt}}(v)$.
@@
We begin with the Bell-nonlocality property of these states. Except for $\rho_{\text{expt}}(v=0)$, which is not even 2-quasiextendible, all other $\rho_{\text{expt}}(v)$ are found to be symmetric $(1,k)$ or $(k,1)$ quasi-extendible iff the corresponding $W^{(d)}(v)$ are. Hence, $\rho_{\text{expt}}(v)$ with $v\geq\frac{1}{2}(1-\frac{2}{k})\neq 0$ cannot violate any Bell inequalities with only $k$ or fewer measurement bases. In fact, despite extensive numerical optimizations, we do not find a Bell-inequality violation by any of the entangled $\rho_{\text{expt}}(v)$; see Section B.1.2. In contrast, using the Horodecki criterion Horodecki et al. (1995), the filtered states $\rho_{\text{expt},f}(v)$ are easily verified to give a maximal CHSH-Bell violation Clauser et al. (1969) of $\mathcal{S}_{\text{\tiny CHSH}}=2.65\pm 0.02,2.46\pm 0.01,2.02\pm 0.02$, and $2.01\pm 0.01$ for $v=0,0.05,0.10$, and $0.15$, respectively, in agreement with the results derived for $W^{(d)}(v)$ Liang et al. (2012). Note, however, that $\rho_{\text{expt},f}(v=0.1)$ and $\rho_{\text{expt},f}(v=0.15)$ violates the CHSH-Bell inequality (with a Bell-local bound of $\mathcal{S}_{\text{\tiny CHSH}}=2$) by less than one standard deviation.
Next, recall from Horodecki et al. (1999) that the usefulness of a quantum state $\rho$ for teleporting an unknown two-qudit state can be confirmed by verifying that its fully entangled fraction:

### v2

@@ Abstract
Introduction.–Ever since its discovery, quantum entanglement Horodecki et al. (2009) has played a fascinating role in shaping Bell (1964) our understanding of the physical world. In the modern era, entanglement takes the role of a powerful resource, enabling quantum computation Jozsa and Linden (2003); Vidal (2003), quantum-enhanced communications Ekert (1991); Bennett and Wiesner (1992), and quantum metrology Tóth and Apellaniz (2014), etc. To take advantage of these possibilities at a large scale, a quantum internet Wehner et al. (2018) where we can readily perform quantum communications between any two points is clearly desirable. Accordingly, the possibility of distributing entanglement in one way or another (see, e.g., Żukowski et al. (1993); Yin et al. (2017)) is surely desirable.
@@
Even with ideal preparation, the quality of entanglement can easily degrade over time, either during the storage or distribution stage. Consequently, much effort has been devoted to understanding how one can recover Bell pairs via entanglement distillation or purification protocols Bennett et al. (1996a); Deutsch et al. (1996). Indeed, for quantum states violating the reduction criterion of separability (RC) Horodecki and Horodecki (1999), iterative single-copy local filtering operations (ScLFs) Gisin (1996) followed by collective local operations Bennett et al. (1996a); Horodecki and Horodecki (1999) distill, with a nonzero probability, quantum states having an arbitrarily high Bell-state fidelity.
@@
Nonetheless, some bipartite entangled states, e.g., those having positive-partial-transposition Peres (1996) (PPT), are undistillable Horodecki et al. (1998). There is even evidence DiVincenzo et al. (2000); Dür et al. (2000) suggesting that some non-PPT (NPPT) quantum states, such as the Werner states Werner (1989), are undistillable. Since all two-qubit entangled states are distillable Horodecki et al. (1997), this intriguing phenomenon of bound entanglement Horodecki et al. (1998) only occurs among higher-dimensional (HD) mixed entangled states. In contrast, pure HD entanglement may offer an advantage over its qubit counterparts in entanglement distribution Steinlechner et al. (2017); Ecker et al. (2019); Hu et al. (2020a) and various information-processing tasks (see, e.g., Liu et al. (2002); Liang et al. (2003); Vértesi et al. (2010)).
@@
Today, it remains unknown Horodecki et al. (2022) whether all NPPT entangled states are distillable. Even among those distillable, many copies may be required Watrous (2004); Fang and Liu (2020, 2022); Regula (2022). Moreover, to distill via the (generalized) recurrence protocol Bennett et al. (1996a); Horodecki and Horodecki (1999), the input state to the protocol must already have a large enough fully-entangled fraction Horodecki et al. (1999) (to the point that it even ensures its usefulness Popescu (1994) for teleportation). Thus, the initial noisy entangled state may have to go through additional ScLFs before being subjected to the collective distillation operations. Even though the latter, more technically challenging, operations have been experimentally realized for two copies Pan et al. (2001, 2003); Zhao et al. (2003); Yamamoto et al. (2003); Walther et al. (2005); Reichle et al. (2006); Kalb et al. (2017) (see also Hu et al. (2021a); Ecker et al. (2021)), we are still far from recovering a near-to-perfect Bell pair via conventional distillation schemes.
@@
In contrast, for the sake of recovering quantum advantage over classical resources, ScLFs often suffice. For instance, they can help Zhao et al. (2017) distinguish different types of entanglement, recover usefulness for secure communication Singh et al. (2021), unveil hidden Pramanik et al. (2019); Nery et al. (2020); Hao et al. (2024); Zhang et al. (2024) quantum steerability Wiseman et al. (2007); Uola et al. (2020), hidden Li et al. (2021) teleportation power Popescu (1994) (see also Masanes (2006)), as well as the (seemingly) hidden Bell-nonlocality of certain two-qubit states Gisin (1996); Kwiat et al. (2001); Wang et al. (2006, 2020) and HD states Popescu (1995); Kumari (2024) (see also Masanes et al. (2008); Liang et al. (2012)). Here, building on the results from Popescu (1995); Liang et al. (2012); Li et al. (2021), we experimentally demonstrate the recovery of various quantum features via the same ScLF for one classic family of HD quantum states—the two-qutrit Werner states. Although tremendous efforts have been devoted to the generation and detection of HD entanglement and their nonlocal properties (see, e.g., Thew et al. (2004); Huber et al. (2010); Schwarz et al. (2016); Martin et al. (2017); Islam et al. (2017); Bavaresco et al. (2018); Zeng et al. (2018); Wang et al. (2018); Hu et al. (2018); Lu et al. (2020); Dada et al. (2011); Hu et al. (2020b, 2021b); Designolle et al. (2021); Qu et al. (2022a, b)), our work gives the first experimental demonstration of the single-copy distillation of quantumness from HD entanglement.
@@
where $\mathbb{I}_{d^{2}}$ is the identity operator acting on $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$, $n_{\pm}^{d}:={\rm tr}\Pi_{\pm}^{d}=\frac{d(d\pm 1)}{2}$ and the swap operator $V$ satisfies $V|\psi\rangle_{A}|\phi\rangle_{B}=|\phi\rangle_{A}|\psi\rangle_{B}$ for all $|\psi\rangle,|\phi\rangle\in\mathbb{C}^{d}$. $W^{(d)}(v)$ are PPT Dür et al. (2000) and separable Werner (1989) if and only if (iff ) $v\geq\frac{1}{2}$. However, entangled Werner states are only known to be distillable when Dür et al. (2000); DiVincenzo et al. (2000) $v<v_{\text{distill}}\equiv\frac{d+1}{4d-2}$. In this regard, a state $\rho$ is said to be distillable Horodecki and Horodecki (2001) if it is $k$-distillable for some integer $k$, i.e., if there exists qubit projectors $\Pi_{A},\Pi_{B}$ such that $\frac{\Pi_{A}\otimes\Pi_{B}\rho^{\otimes k}\Pi_{A}\otimes\Pi_{B}}{{\rm tr}(\Pi_{A}\otimes\Pi_{B}\rho^{\otimes k}\Pi_{A}\otimes\Pi_{B})}$ is entangled. To determine the existence of NPPT bound entangled states Horodecki et al. (2022), it suffices Dür et al. (2000) to consider $W^{(d)}(v)$ for $d>2$.
@@
HD Werner states are not only important because of their role in the NPPT distillability problem but are also peculiar for being useless in various tasks. Indeed, all HD $W^{(d)}(v)$ satisfy RC Horodecki and Horodecki (1999). Consequently, even those that are distillable cannot be fed directly to the generalized recurrence protocol Horodecki and Horodecki (1999) for entanglement distillation. Their compliance with RC also renders them useless for quantum teleportation Horodecki and Horodecki (1999) and dense coding Bruß et al. (2004). However, as shown in Li et al. (2021), ScLF in the form of qubit projection can activate their usefulness for teleportation whenever $v<v_{\text{distill}}$.
@@
Furthermore, Werner states with $v\geq v_{\text{steer}}\equiv\frac{d+1}{2d^{2}}$ are the first known examples of entangled quantum states not violating any Bell Werner (1989) or steering Wiseman et al. (2007) inequality with projective measurements (see Fig. 1). However, after a successful qubit ScLF Popescu (1995), $W^{(d\geq 5)}(v_{\text{steer}})$ does violate the Clauser-Horne-Shimony-Holt (CHSH) Bell inequality Clauser et al. (1969). More generally, $W^{(d)}(v)$ admit (Johnson and Viola, 2013, Theorem 6) a symmetric $(1,k)$ or $(k,1)$-extension Terhal et al. (2003); Doherty et al. (2002) iff $v\geq v_{\text{Sym}}\equiv\frac{1}{2}\left(1-\frac{d-1}{k}\right)$. In other words, there exists a $(1+k)$-partite qudit state $\rho$ that recovers $W^{(d)}(v)$ after tracing out any $(k-1)$ copies of Alice’s (Bob’s) subsystems. It then follows Terhal et al. (2003) that $W^{(d)}(v)$ for $v\geq v_{\text{Sym}}$ cannot violate any Bell inequality with $k$ or fewer measurement settings for Alice (Bob) Not . In particular, since $v_{\text{Sym}}$ vanishes when $k=d-1$, $W^{(d)}(v)$ satisfy Terhal et al. (2003) all Bell inequalities with $d-1$ or fewer measurement bases at one side. For example, all $W^{(3)}(v)$ cannot violate any two-or-fewer-setting Bell inequality on one side. Again, a qubit ScLF reveals Liang et al. (2012) the Bell-nonlocality (possibly) hidden in some of these states, which we demonstrate experimentally in this work.
@@
How does a qubit ScLF enable the distillation of many of these useful properties? We can gain some intuition into this by first noting that $\Pi_{+}^{d}+\Pi_{-}^{d}=\mathbb{I}_{d^{2}}$ and Popescu (1995) $\Pi_{-}^{d}=\sum_{i<j}\varrho_{0}^{i,j}$ where $\varrho_{0}^{i,j}=|\Psi^{-}_{i,j}\rangle\!\langle\Psi^{-}_{i,j}|$ is a singlet projector defined on the two-qubit subspace $\mathcal{H}_{i,j}:=\text{span}\{|k\ell\rangle:=|k\rangle|\ell\rangle\}_{k,\ell=i,j}$. Then, by defining $p:=\frac{1}{d}$, $q:=1-\frac{2d}{d+1}v$, and noting that $n_{-}^{d}$ is the binomial coefficient of $d$ chooses 2, we can rewrite $W^{(d)}(v)$ as a uniform mixture of two-qubit states $\varrho^{i,j}$:
@@
where $\varrho_{1}^{i,j}=\frac{|ii\rangle\!\langle ii|+|jj\rangle\!\langle jj|}{2}$, $\varrho_{2}^{i,j}=\frac{|ij\rangle\!\langle ij|+|ji\rangle\!\langle ji|}{2}$. In this decomposition, only the two-qubit singlet projectors $\{\varrho_{0}^{i,j}\}_{i,j}$ contribute to the entanglement of $W^{(d)}(v)$. The act of a qubit projection, say, $\Pi_{2}:=|0\rangle\!\langle 0|+|1\rangle\!\langle 1|$, on both sides then keeps the maximally entangled $\varrho_{0}^{0,1}$ and the separable $\frac{\mathbb{I}_{4}}{4}$ with some probability while removing the components in all other subspaces, thereby effectively “purifying” the entanglement contained in $W^{(d)}(v)$. From Eq. 2a, we see that upon a successful qubit projection, since $\Pi_{2}\otimes\Pi_{2}W^{(d)}(v)\Pi_{2}\otimes\Pi_{2}\propto W^{(d=2)}(v^{\prime})$, we obtain a two-qubit Werner state with a symmetric weight $v^{\prime}=\frac{3(d-1)v}{(d+1)(1-v)+3(d-1)v}$. Note that $v^{\prime}\geq v$ for all $d>2$.
@@
Experimental Werner states and filtering operation.–The decomposition of Eq. 2 not only sheds light on the relevance of a qubit ScLF but also provides a systematic way for preparing an entangled $W^{(d)}(v)$ for an arbitrary $d$ via generating and mixing the various two-qubit states $\varrho^{i,j}(q)$ residing in $\mathcal{H}_{i,j}$. To this end, we start by preparing photons pairs maximally entangled in the polarization degree of freedom (DOF) $|\Psi^{+}\rangle_{\text{AB}}=\frac{1}{\sqrt{2}}(|H\rangle|V\rangle+|V\rangle|H\rangle)_{\text{AB}}$, where $H$ and $V$ denote, respectively, horizontal and vertical polarization. As shown in Fig. 2a, the entangled photons are generated via the spontaneous parameter down conversion process (SPDC) from a periodically-poled potassium titanyl phosphate (PPKTP) crystal in a Sagnac interferometer, which is bidirectionally pumped by a laser with central wavelength at 405 nm. By setting a HWP at 0 on photon B, $|\Psi^{+}\rangle_{\text{AB}}$ is transformed into $\varrho_{0}^{H,V}=|\Psi^{-}\rangle_{\text{AB}}\langle\Psi^{-}|$. Then, we use two BSs, two PBSs and two tunable LAAs (with transmittance of $q$ and $1-q$, respectively) to partially decohere $\varrho_{0}^{H,V}$ and prepare the mixed state $\varrho^{H,V}(q)$ in Eq. 2 by randomly rotating a HWP between $45^{\circ}$ and $0^{\circ}$. Finally, we obtain the $d$-dimensional Werner states $W^{(d)}(v)$ from $\varrho^{H,V}(q)$ with a qubit-to-qudit mapping, where the qudit is encoded in the path DOF $\{|l_{k}\rangle\}$ and the mapping is achieved with two optical networks each consisting of $d-1$ beam displacers (BDs) and a series of HWPs on photon A and B. For $d=3$ in this work, we use two BDs and two HWPs at each side (as shown in Fig. 2a) to prepare eleven $W^{(d=3)}(v)$ with $v\in[0,0.5]$ in steps of 0.05. See Section A.2 for further details.
@@
ScLF and distillation of quantumness.–With our encoding of the quantum information, a qubit filtering—relevant for distilling the various quantum features alluded to above—is straightforward. To keep only contributions to the two lower optical paths $\{|l_{1}\rangle,|l_{2}\rangle\}$, we block the upper optical path $|l_{0}\rangle$ (Fig. 2b). After ScLF, the two-qubit state $\rho_{\text{expt},f}(v)$ encoded in $\{|l_{1}\rangle,|l_{2}\rangle\}$ is mapped back to the polarization DOF and reconstructed using a HWP, a QWP and a PBS—their fidelity ranges roughly from 0.934 to 0.995. In the following, we compare the properties of $\rho_{\text{expt}}(v)$ and $\rho_{\text{expt},f}(v)$ to illustrate how ScLF has enabled the recovery of quantum features missing in $\rho_{\text{expt}}(v)$.
@@
We begin with their Bell-nonlocality property. Except for $\rho_{\text{expt}}(v=0)$, which is not even 2-quasiextendible, all other $\rho_{\text{expt}}(v)$ are symmetric $(1,k)$ or $(k,1)$ quasi-extendible iff the corresponding $W^{(d)}(v)$ are. Hence, $\rho_{\text{expt}}(v)$ with $v\geq\frac{1}{2}(1-\frac{2}{k})\neq 0$ cannot violate any Bell inequalities with $k$ or fewer measurement bases at one side. In fact, despite extensive numerical optimizations, we do not find any Bell-inequality violation by $\rho_{\text{expt}}(v)$; see Section B.2. In contrast, using the Horodecki criterion Horodecki et al. (1995), the filtered states $\rho_{\text{expt},f}(v)$ give a maximal CHSH-Bell value Clauser et al. (1969) of $\mathcal{S}_{\text{\tiny CHSH}}=2.65\pm 0.02,2.46\pm 0.01,2.02\pm 0.02$, and $2.01\pm 0.01$ for $v=0,0.05,0.10$, and $0.15$, respectively, in agreement with the results derived for $W^{(d)}(v)$ Liang et al. (2012). Note, however, that $\rho_{\text{expt},f}(v=0.1)$ and $\rho_{\text{expt},f}(v=0.15)$ violates the CHSH-Bell inequality ($\mathcal{S}_{\text{\tiny CHSH}}\leq 2$) by less than one standard deviation.
Next, recall from Horodecki et al. (1999) that a quantum state $\rho$ is useful for teleporting an unknown two-qudit state if its fully-entangled fraction: